10. If a is a unit vector which is a function of a scalar prove that da/dt, in general a and da/dt are at right angles.
1
Expert's answer
2020-01-21T03:32:44-0500

Let's us assume unit vector as "a=cos \ \theta \ \hat { i }\ +sin\ \theta\ \hat{j}"

calculating the "da/dt\ =-(sin\ \theta)\ \hat{i}\ +cos\ \theta\ \hat{j}"

so the vector "a" and "da\/dt" are perpendicular if their dot product are zero

"a.da\/dt=0"

hence vector "a" and "da\/dt" are perpendicular.